Express 2/3 degrees in minutes
Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(5404)

Express 2/3 degrees in minutes
.
Since there are 60 minutes per 360 deg
we multiply 2/3 by 60/360:
.
2/3 * 60/360
2/3 * 1/6
2/18
1/9 minutes

Answer by Alan3354(21599)

If you mean minutes of arc, there are 60' per degree
2/3º = (2/3)*60' = 40'